### v2

@@ Abstract
This paper presents a secure and versatile sampling-based verification protocol, Proof of Sampling (PoSP) protocol, suitable for a wide range of decentralized applications. Our protocol has a unique Nash Equilibrium in pure strategies, which compels rational participants to act honestly, thus fortifying the network’s integrity. This can be achieved with manageable computational overhead. When applied to decentralized inference for AI applications, we design spML based on PoSP protocol, which ingeniously amalgamates the strengths of optimistic fraud proof and zero knowledge proof based approaches, the foremost approaches in the domain at present. Moreover, the PoSP protocol can be effectively utilized for designing verification mechanisms within Actively Validated Services (AVS) in EigenLayer and Layer 2 solutions, further broadening its applicability. This innovative approach not only enhances the security and efficiency of decentralized systems but also paves the way for a new generation of scalable and reliable decentralized applications.
@@ 1 Introduction
In the development of decentralized protocols, it is customary to presuppose that honest nodes will adhere to the established protocol. Take, for instance, optimistic rollup [7] is a scaling solution for blockchain that aims to increase blockchain’s transaction throughput. In this approach, a designated rollup validator processes transactions off-chain and posts the results on the blockchain. By default, the transactions are assumed to be executed correctly. However, if other validators detect incorrect transactions, they can submit fraud proofs on-chain to challenge the validator who submitted the false results. The security of optimistic rollup relies on the critical assumption that at least one rollup validator is honest. If all validators are dishonest, the on-chain transactions could be fraudulent. To incentivize honest behavior, an economic reward system is introduced, making it more profitable for rational nodes to act honestly. However, under this assumption, [14] observed that in most existing decentralized systems, the equilibrium state corresponds to a mixed strategy Nash Equilibrium, which means that the optimal strategy for each validator is to cheat a certain percentage of the time! This implies a non-negligible probability of dishonest behavior that is never detected and introduces significant security vulnerabilities. Furthermore, [14] also proposed an alternate framework for optimistic rollups that mandates the scrutiny of the system by all nodes. However, such a design introduces duplicated work from all nodes and contravenes the fundamental principle of scalability, undermining the initial objective of Layer 2 design.
In this paper, we introduce the Proof of Sampling (PoSP) protocol designed to address these challenges effectively, under certain foundational assumptions. The PoSP protocol is applicable across a broad range of decentralized systems. In this paper, we introduce its integration into decentralized AI inference platforms. Notably, our system achieves a unique Nash Equilibrium in pure strategies, wherein each node acts honestly, an principle that is highlighted in [9] as ideal for the design of secure decentralized systems. This signifies that even when each node’s strategy is directed towards maximizing its individual profit, the overall system maintains optimal security.
@@ 1.1 Main Contribution
The main contributions of this paper are twofold. First, we propose the PoSP protocol, a robust protocol for a wide range of decentralized applications. PoSP protocol leverages a unique Nash Equilibrium in pure strategies to incentivize rational actors within the network to maintain honest behavior, thereby significantly reinforcing the network’s security and integrity. Second, we design spML, a sampling-based verification mechanism for decentralized AI inference network, using PoSP protocol. SpML synthesizes the advantages of both optimistic fraud proof and zero knowledge proof based approaches (opML and zkML), incorporating the high scalability and simplicity of opML with the enhanced security features of zkML. SpML stands out for its strategic balance of computational efficiency and security, positioning it as a superior alternative within the decentralized AI inference landscape.
In sum, our work not only introduces PoSP as a foundational protocol for secure and efficient verification in decentralized systems but also presents the development of spML, a balanced approach in decentralized AI, derived from the PoSP protocol. These innovations significantly advance the field, fostering a new generation of decentralized applications.
1.2 Design Principles
PoSP enhances decentralized system security by initially selecting an asserter who submits results without knowing which validators might review them. If challenged, $n$ validators are randomly selected, and only then do they learn each other’s identities. This approach prevents collusion at the outset and limits free-riding by concealing the results until all verifications are complete, ensuring independent validation by each participant.
In the case where all nodes agree on the outcome, the system accepts this consensus, ensuring that the protocol functions correctly. However, if there are discrepancies, indicating dishonesty or error, those nodes are penalized. This mechanism ensures that acting honestly is not only the most straightforward but also the most beneficial strategy for all participants, thus maintaining the integrity and security of the system.
1.3 Literature Review
@@
Employing game theory for protocol analysis, rational and economic theories have been extensively applied to explore various blockchain configurations, including Byzantine Fault Tolerance as discussed in [1, 6], sharding strategies [15], proof-of-work systems [2, 5, 10, 13], proof-of-stake systems [3, 16, 17], secure outsourced computation [20], and Layer 2 solutions [11, 12, 14, 18]. The use of slashing mechanisms, analyzed in [9], enhances network security and ideally aims to align node incentives in such a way that following the protocol-prescribed strategy always yields the most benefit, reflecting a unique Nash Equilibrium in pure strategies. A comprehensive review [21] elaborates on the diverse applications of game theory across different blockchain scaling solutions. This analytical approach underscores the utility of game theory in dissecting and enhancing the security and scalability mechanisms of blockchain technology.
@@
The decentralized AI domain has significantly benefited from innovations like opML, as detailed by [3] and [19], which enhances machine learning (ML) on the blockchain by facilitating the efficient processing of intricate models, alongside the introduction of Zero-Knowledge Machine Learning (zkML) by [22, 23]. These advancements in blockchain AI, opML and zkML, address scalability, security, and efficiency with distinct trade-offs. OpML enhances scalability and efficiency, yet its security may not be as robust. ZkML offers strong security through zero-knowledge proofs, yet faces challenges in scalability and efficiency due to high computational overhead.
@@
The rest of this paper is organized as follows. In Section 2, we introduce PoSP protocol and prove that it has a unique Nash Equilibrium in pure strategies. PoSP protocol can be applied in all kinds of verification scenarios. In Section 3, we show a possible way to implement PoSP protocol in real world applications as an example. We design a sampling-based verification mechanism (spML), which is implemented by PoSP protocol, within a decentralized AI inference network, detailing the protocol design and its security validation. In Section 4, we conclude the paper and discuss the future extensions of our work.
2 Model
2.1 PoSP Protocol
In this section, we propose the PoSP protocol.
@@ 2.1 PoSP Protocol
    A node is randomly selected from the network to serve as an asserter. This asserter calculates a value $f(x)$, with both the function $f(\cdot)$ and the input $x$ being well-known to the network, and outputs the result.
@@
    With a predetermined probability $p$, a challenge mechanism protocol is triggered. If the challenge mechanism is not triggered, this round concludes, and the asserter is awarded a reward denoted by $R_{A}$.
@@
    If the challenge mechanism is triggered, $n$ validators are randomly selected from the network, where $n\geq 1$ is a predetermined integer parameter. Each validator, denoted as validator $i$, independently computes $f(x)$ and outputs the result.
@@
    If all of the results from the asserter and the valitors match, the result is deemed valid and accepted. The asserter receives $R_{A}$ and each validator receives $R_{V}/n$. Otherwise, an arbitration process is initiated to determine the honesty of the asserter and each of the $n$ validators. If the asserter is proven honest, it receives a reward of $R_{A}$. Otherwise, it is penalized with $S$. Each honest validator, out of the $m$ honest ones from a total of $n$ validators, receives $R_{V}/m$ if $m\geq 1$, while each dishonest validator is penalized with $S$.
@@ 2.2 Assumptions and Analysis
Then we show that, under certain conditions, PoSP protocol has a unique Nash Equilibrium in pure strategies that all nodes act honestly. This ensures there is no economic incentive for nodes to perform malicious actions. To analyze this protocol, we additionally define
  • $C$: computational cost for computing $f(x)$
@@
  • $U_{2}$: maximum profit that the asserter can gain if the challenge mechanism is triggered and he colludes with all validators
@@
We assume that if the validator does not collude with the asserter, their messages will never match with each other unless they both act honestly in the network.
@@
We also assume that $R_{A}-C>-S$ and $R_{V}/n-C>-S$, which means that it is better for each node to receive the reward than be slashed.
Property 1
Under Assumption 1 and Assumption 2, if one validator does not collude with the asserter, his dominant strategy is to act honestly.
Strategy Asserter Fraud Asserter Not Fraud
Validator Honest $R_{V}/m-C$ $R_{V}/m-C$
Validator Dishonest $-S$ $-S$
The table above gives a game in a bimatrix format. The first row is the behavior of the asserter, and the first column is the behavior of the validator, who does not collude with the asserter. The value in the table is the payoff of the validator. In this case, Property 1 is straightforward by the table.
Strategy Detected Undetected
Assserter Fraud $-S$ $U_{1}$ or $U_{2}$
Asserter Not Fraud $R_{A}-C$ $R_{A}-C$
The table above shows the payoff of the asserter. The first row represents whether his behavior will be detected by the protocol, and the first column denotes the behavior of the asserter.
Assumption 3
We assume that the fraction of the nodes in the network that the asserter controls is $r$, which means at most a fraction of $r$ nodes are able to collude with each other.
@@
Under Assumption 1, Assumption 2 and Assumption 3, the asserter has a dominant strategy to act honestly if and only if
$R_{A}+pS-(1-p)U_{1}-C+(1-r^{n}-(1-r)^{n})pC>pr^{n}(U_{2}+S-R_{V}).$
@@
Corollary 1
If
$R_{A}+pS-(1-p)U_{1}-C>pr^{n}(U_{2}+S-R_{V}),$
the asserter has a dominant strategy to act honestly.
Proof
We can prove this Corollary by the inequality of $1-r^{n}-(1-r)^{n}\geq 0$ and Theorem 2.1.
@@ 3 Application to Decentralized AI Inference Network Design
With the rapid growth in AI, centralized AI servers and corporations are beginning to encounter significant limitations. For instance, it is nearly impractical for a limited number of centralized servers to fulfill the burgeoning global demand for AI inference. Consequently, the concept of a decentralized AI inference network has gained popularity in recent times. Such a network leverages the computational power of a wide array of individual providers who contribute to the AI server pool in a permissionless manner. A well-designed decentralized AI inference network is able to balance the supply with the demand for AI inference capabilities.
In a decentralized network, simply executing a model and trusting its output is insufficient. Consider a scenario where the network is requested to analyze a complex governance issue using a sophisticated model, such as Llama2–70B. There arises a critical question: how can one be certain that the network is not employing a less capable model, like Llama2–13B, thereby delivering subpar analysis while surreptitiously profiting from the discrepancy?
In a centralized setting, one might place their trust in established companies like OpenAI, largely because such entities have a reputation to uphold. Moreover, the quality of a Language Learning Model tends to be self-evident to some degree. However, in a decentralized environment, trust cannot be taken for granted—it must be established through rigorous verification.
@@
SpML is ingeniously designed to excel in the verification process, thereby ensuring the integrity of the system with a negligible increase in computational overhead for security purposes. It stands out as a particularly rational solution when compared to other methods that have been proposed. The following sections will detail the application of PoSP protocol in establishing a robust decentralized AI inference network.
3.1 System Architecture

### v3

@@ Abstract
This paper introduces the Proof of Sampling (PoSP) protocol, a Nash Equilibrium-based verification mechanism, and its application to decentralized machine learning inference through spML. Our protocol has a pure strategy Nash Equilibrium, compelling rational participants to act honestly. It economically disincentivizes dishonest behavior, making it costly for participants to compromise the network’s integrity.
In our spML protocol, we apply PoSP to decentralized inference for AI applications via a novel cryptographic protocol. The resulting protocol is much more efficient than zero knowledge proof based approaches. Moreover, we anticipate that the PoSP protocol could be effectively utilized for designing verification mechanisms within Actively Validated Services (AVS) in restaking solutions.
We further expect that the PoSP protocol could be applied to a variety of other decentralized applications. Our approach enhances the reliability and efficiency of decentralized systems, paving the way for a new generation of decentralized applications.
@@ 1 Introduction
In the development of decentralized protocols, it is customary to assume that honest nodes will adhere to the established protocol. Take, for example, optimistic rollups [14] as a scaling solution for blockchain that aims to increase blockchain’s transaction throughput. In this approach, a designated rollup validator processes transactions off-chain and posts the results on the blockchain. By default, the transactions are assumed to be executed correctly. However, if other validators detect incorrect transactions, they can submit fraud proofs on-chain to challenge the validator who submitted the false results. The security of optimistic rollup relies on the critical assumption that at least one rollup validator is honest and validates the transactions. If all validators are dishonest, the on-chain transactions could be fraudulent. To incentivize honest behavior, an economic reward system is introduced, making it more profitable for rational nodes to act honestly. However, under this assumption, Mamageishvil and Felten[25] observed that in most existing decentralized systems, the equilibrium state corresponds to a mixed strategy Nash Equilibrium, which means that the optimal strategy for each validator is to cheat and not validate a certain percentage of the time. In addition, [25] also proposed an alternate framework for optimistic rollups that mandates the scrutiny of the system by all nodes. However, such a design introduces duplicated work from all nodes and contravenes the fundamental principle of scalability, undermining the initial objective of Layer 2 design.
In this paper, we introduce the Proof of Sampling (PoSP) protocol designed to address these challenges effectively, under certain foundational assumptions. We expect that the PoSP protocol is applicable across a broad range of decentralized systems. In this paper, we introduce its application to decentralized AI inference platforms. Notably, our system achieves a pure strategy Nash Equilibrium, wherein each node outputs the correct result, a principle that is highlighted by [16] as ideal for the design of secure decentralized systems. This signifies that when each node’s strategy is directed towards maximizing its individual profit, the overall system maintains security.
@@ 1.1 Main Contribution
We made two contributions in this paper:
  • First, we propose Proof of Sampling (PoSP), a pure strategy Nash Equilibrium protocol to incentivize rational actors within the network to return the correct output, thereby strengthening the security and integrity of the decentralized protocol.
  • Based on PoSP, we design spML, a concrete instantiation of the PoSP protocol for decentralized machine learning inference. Assuming all nodes act rationally, spML ensures that each node is incentivized to act correctly under some reasonable trust assumptions.
1.2 Related Work
@@
Game Theory for Protocol Analysis. Employing game theory for protocol analysis, rational and economic theories have been extensively applied to explore various blockchain configurations, including Byzantine Fault Tolerance as discussed in [1, 12], sharding strategies [26], proof-of-work systems [3, 10, 17, 24], proof-of-stake systems [5, 31, 32], secure outsourced computation [35], and Layer 2 solutions [13, 19, 20, 21, 25, 33]. The use of slashing mechanisms, analyzed in [16], enhances network security and ideally aims to align node incentives in such a way that following the protocol-prescribed strategy always yields the most benefit, reflecting a unique Nash Equilibrium in pure strategies. A comprehensive review [36] elaborates on the diverse applications of game theory across different blockchain scaling solutions. This analytical approach underscores the utility of game theory in dissecting and enhancing the security and scalability mechanisms of blockchain technology.
@@
Decentralized AI Inference. The decentralized AI domain has significantly benefited from innovations like opML, as detailed by[7], which enhances machine learning (ML) on the blockchain by facilitating the efficient processing of intricate models, alongside the introduction of Zero-Knowledge Machine Learning (zkML) by [2, 15, 22, 29, 34, 37, 38, 39]. These advancements in blockchain AI, opML and zkML, address scalability, security, and efficiency with distinct trade-offs. OpML enhances scalability and efficiency, yet its security may not be as robust. ZkML offers strong security through zero-knowledge proofs, yet faces challenges in scalability and efficiency due to high computational overhead.
Game Design with Pure Strategy Nash Equilibrium. Among the most relevant studies we reviewed, the following papers stand out for their design of game-theoretic models that achieve a unique Nash Equilibrium in pure strategies, where each node is incentivized to act honestly. This core design criterion is fundamental to our research as it aligns closely with our objectives of ensuring integrity and reliability in decentralized network environments.
Outsourced computation within game theoretic frameworks, as discussed in studies [4, 18], commonly involves a trusted third party to resolve disputes and deter collusion among coalitions. In our paper, we can use the arbitration protocol to ensure integrity without relying on a trusted third party.
Nix and Kantarcioglu [28] implements a system where the integrity of outsourced data is maintained through a verification game involving two non-colluding cloud providers. In this model, the data owner begins by sending a query to one of the cloud providers. With a predefined probability, this query is also sent to a second provider for computation. The results from both providers are then compared to detect any discrepancies, effectively using one provider to challenge the results of the other. This approach assumes that the nodes do not collude and that there is a negligible probability of their results matching if both are incorrect. In contrast, our framework does not rely on these assumptions. We accommodate the possibility of collusion and does not rely on the assumption that non-colluding parties will always produce differing results if both are dishonest. This makes our protocol more realistic and applicable to real-world scenarios.
Dong et al.[8] discusses smart counter-collusion contracts for verifiable cloud computing; it utilizes game theory to foster distrust among potential colluders, effectively discouraging collusion by incentivizing betrayal among partners, making it economically disadvantageous and risky. While their model relies on the rational behavior of physically isolated nodes and assumes the presence of a trusted third party for dispute resolution, our approach offers a broader application by accommodating the possibility of irrational collusion and the control of multiple nodes by a single entity. Instead of relying on a centralized trusted third party, our model distributes trust to ensure integrity and resolve disputes, better aligning with the decentralized nature of blockchain environments and enhancing system robustness against a wider range of threats. This makes our framework more adaptable to real-world decentralized settings.
Lu et al. [23] presents a mechanism for decentralized blockchain systems without a trusted third party. In their model, if the results from multiple nodes differ, all nodes are penalized to prevent collusion and promote individual integrity. This approach addresses the issue that, in large networks, one party might control a significant number of nodes $n-1$, yet $n$ nodes are required to participate in each round. In contrast, our approach is more refined; it considers scenarios where a certain fraction of nodes might be controlled by one entity. In most cases, such as the application for decentralized AI inference network, we only employ a two-node system per round, randomly selecting a validator unknown to the asserter until after submission, which is a key point for our design. Additionally, our system uses blockchain-based arbitration to resolve discrepancies, ensuring that honest nodes are not unjustly penalized, thus maintaining fairness and enhancing reliability without relying on a trusted third party.
Pham et al. [30] outlines the use of a trusted third party for auditing outsourced computations, where discrepancies between nodes’ results trigger an audit, and random checks ensure integrity even when results match. In contrast, our work eliminates the need for a trusted auditor by leveraging distributed trust to arbitrate disputes only when results differ, enhancing decentralization and reducing computational overhead. We specifically address potential node collusion, assuming a scenario where up to a certain fraction of the network might collude or be controlled by a single party. Our paper not only simplifies the verification process by relying on result comparison and arbitration, which never happens if every node is rational, rather than correctness checks with a certain probability but also strategically mitigates collusion risks by maintaining validator anonymity until asserter’s submissions are finalized.
@@
The rest of this paper is organized as follows. In Section 2, we introduce PoSP protocol and prove that it has a unique Nash Equilibrium in pure strategies under certain conditions. In Section 3, we show a possible way to implement PoSP protocol in real world applications as an example. We design a sampling-based verification mechanism (spML), which is implemented by PoSP protocol, within a decentralized AI inference network, detailing the protocol design and its security validation. In Section 4, we conclude the paper and discuss the future extensions of our work.
2 The PoSP Protocol
The PoSP (Proof of Sampling) protocol is developed from a game theoretic perspective, focusing on the strategic interactions within a multi-agent system. This approach emphasizes the decision-making processes and the incentive structures that are essential for ensuring that the actions of all nodes, whether potentially Byzantine or inherently honest, align with the network’s objectives. Specifically, the PoSP protocol is structured to achieve a unique Nash Equilibrium in pure strategies, where every Byzantine participant is incentivized to act honestly, thereby enhancing the reliability and integrity of the system. While this section lays the theoretical groundwork, the specific application and implementation details are context-dependent. In the subsequent section, we will explore the application of PoSP in decentralized AI inference, specifically in a scenario termed spML, based on our PoSP protocol. Detailed implementation of spML will be discussed there.
2.1 Protocol Design
In this section, we propose the PoSP protocol. Consider a system with $N$ nodes. We assume that $f$ is a deterministic function, and $x$ is an input. An amount $B$ is paid for the following transaction computing $f$ on $x$. Let $R_{A}$ and $R_{V}$ be two positive numbers such that $R_{A}+R_{V}<B$.
@@ 2.1 Protocol Design
    A node selected from the network serves as an asserter. This asserter calculates a value $f(x)$, with both the function $f(\cdot)$ and the input $x$ being well-known to the network, and outputs the result.
@@
    With a predetermined probability $p$, a challenge protocol is triggered. If the challenge protocol is not triggered, this round concludes, and the asserter is awarded a reward denoted by $R_{A}$.
@@
    If the challenge protocol is triggered, $n$ validators are randomly selected from the network of $N$ nodes, where $n\geq 1$ is a predetermined integer parameter. Each validator, denoted as validator $i$, independently computes $f(x)$ and outputs the result.
@@
    If all results from the asserter and the validators match, the result is deemed valid and accepted. The asserter receives $R_{A}$ and each validator receives $R_{V}/n$. Otherwise, an arbitration protocol is initiated to determine the correctness of the asserter’s result and that of each of the $n$ validators. We assume that the arbitration process is always accurate. If the asserter’s result is upheld by the arbitration protocol, it receives a reward of $R_{A}$; if not, it is penalized with $S$. Similarly, each validator deemed incorrect will also be penalized with $S$. Let $S_{\mathsf{total}}$ represent the total amount slashed from dishonest validators and the asserter. Each validator deemed honest, out of the $m$ deemed-honest ones from a total of $n$ validators, receives $R_{V}/m+S_{\mathsf{total}}/m$ if $m\geq 1$.
In the PoSP protocol, the combined values of all rewards and penalties are carefully designed to ensure that the amount paid out never exceeds the amount collected. Excess tokens, arising from scenarios like untriggered challenges or accumulated penalties, are burned or go to the protocol reserve. This cautious approach also helps prevent manipulative behavior and maintains the system’s integrity.
@@ 2.2 Assumptions and Analysis
Then we show that, under certain conditions, PoSP protocol has a unique Nash Equilibrium in pure strategies 2.1 that all nodes output the correct result. This ensures there is no economic incentive for nodes to produce erroneous results. To analyze this protocol, we additionally define
  • $C$: the computational cost for a reference implementation of $f(x)$, established to ensure that all nodes allowed in the network can accurately compute the function within this cost
@@
  • $U_{2}$: maximum profit that the asserter can gain if the challenge mechanism is triggered and he controls all the validator nodes
@@
We assume that $S>nC$, $R_{A}-C>-S$ and $R_{V}/n-C>-S$.
These inequalities imply that, it is more beneficial for each node to receive the reward than to be penalized. Then we show that given Assumption 1, the Byzantine validator who does not collude with the asserter has a dominant strategy to output the correct result.
Strategy Asserter Correct Asserter Incorrect
Validator Correct $\geq R_{V}/n-C$ $\geq R_{V}/n-C+S/n$
Validator Incorrect $-S$ $\leq R_{V}/n$
Table 1: Payoff for Validator in a Non-Collusive Scenario
Table 1 gives a game in a bimatrix format. The first row is the correctness of the asserter’s outcome, and the first column is the correctness of the validator’s outcome, who does not collude with the asserter. The value in the table is the payoff of the validator. Then, the following property is straightforward.
Property 1
Under Assumption 1, if one validator does not collude with the asserter, its dominant strategy is to output the correct result.
@@
We assume that at most a fraction $r$ of the total nodes in the network are Byzantine (that is, might deviate from the correct execution of $f$), which also implies that the fraction of the nodes in the network that the Byzantine asserter controls is at most $r$.
@@
Under Assumption 1 and Assumption 2, each asserter has a dominant strategy to act honestly and output the correct result if
$R_{A}+pS-(1-p)U_{1}-C>pr^{n}\left(U_{2}+S\right).$
This also means our system has a unique Nash Equilibrium in pure strategies, where each node behaves honestly and reliably outputs the correct result.
@@ 3 Application to Decentralized AI Inference Network Design
Decentralized AI inference network has gained popularity in recent times due to rapid growth of AI and strong demand from the industry. Such a network leverages the computational power of a wide array of individual providers who contribute to the AI server pool in a permissionless manner. A well-designed decentralized AI inference network is able to balance the supply with the demand for AI inference capabilities.
In a decentralized setting, AI inference services are performed by these distributed nodes. However, these nodes are not guaranteed to behave honestly, so simply executing a model and trusting its output is insufficient. For instance, if a user wants to perform AI inference using a powerful model like LLaMA3-70B, an execution node might choose to use a less capable model like LLaMA3-7B to save computational power. Therefore, it is crucial to design additional mechanisms that incentivize honest execution and penalize nodes that act dishonestly.
@@
SpML is designed to verify the integrity of the system with a minimal increase in computational overhead for security purposes. The following sections will detail the application of the PoSP protocol in establishing a robust decentralized AI inference network.
3.1 Design Principles & Assumptions
Deterministic ML Execution. One critical assumption that PoSP makes, which spML inherits, is that the function $f$ must be a determinstic function. Achieving deterministic execution in ML is notoriously diffcult due to the inherent inconsistent nature of floating point arithmetics, as outlined in [40]. For example, multiplying two floating point numbers on different software/hardware configuration might render different reuslts.